# Symbolic Constraints in Polyhedral Enclosure and Tetrahedral Decomposition in Genus-0 Polyhedra

### v5

@@ 5.4 Count of Valid Internal Configurations
5.5 External Constraints on Internal Structure
While for polyhedra of minimal vertex count, surface flatness $S$ typically correlates with increased constraint on internal decomposition, restricting the range of valid $(T,N_{i},S_{i})$ configurations, this relationship becomes unreliable at higher vertex counts. In essence, internal complexity is not solely determined by how flat a polyhedron’s surface is, but also by how that flatness is distributed.
For $S=0$, the full tetrahedral range is always accessible, as no surface constraint exists.
At higher values of $S$, internal structure is expected to compress. However, I observe that when flatness is **distributed across many external faces**, the volume cannot be enclosed efficiently with shallow tetrahedral configurations alone. This leads to the appearance of fully internal tetrahedra, a phenomenon I refer to as lettucing. In these cases, internal layering is necessary to fill the bulk.
By contrast, when flatness is concentrated on one or two faces, for example, in a polyhedron with a large polygonal base or cap, the resulting internal tetrahedra tend to form a single radial layer, each anchored to the large flat surface. These configurations are shallow: all tetrahedra have at least one face or edge exposed.
These behaviors become increasingly common and complex as $V$ increases, due to the geometric and combinatorial expansion of allowable configurations. Thus, while heuristic rules can be proposed for internal structure at a given $V$, e.g., $4\leq V\leq 8$, such rules are often misleading, as they obscure emergent phenomena like volumetric layering that only manifest at higher vertex counts.
@@ 5.6 Example for $V=8$
$S$ (External) $T$ (Tetrahedra) $N_{i}$ (Triangles) $S_{i}$ (Segments)
0 5 4 0
0 6 6 1
0 7 8 2
0 8 10 3
1 5 4 0
1 6 6 1
1 7 8 2
2 5 4 0
2 6 6 1
3 5 4 0
4 5 4 0
5 5 4 0
6 5 4 0
From these configurations, the following heuristic rules might be proposed:
  • When $S=0$: All V4V-4 configurations are valid.
  • When $S=1$: The configuration with $T_{\max}$, $N_{i\,\max}$, and $S_{i\,\max}$ is excluded.
  • When $S=2$: The top 2 configurations are excluded.
  • When $S=k$: Configurations with $T\in[T_{\max}-(k-1),T_{\max}]$ are excluded.
  • When $S\geq V-5$: The number of valid configurations becomes narrow, and one might assume it reduces to the minimal configuration $(T_{\min},N_{i\min},S_{i\min})$. However, this assumption fails immediately with the inclusion of one additional vertex.
Consider the following two polyhedra, both with $V=9$ and $S=5$:
  • An octagon capped by eight triangles (a domed structure): flatness is concentrated on one side. It requires only $T=6$ tetrahedra, with $S_{i}=0$ and $N_{i}=5$.
  • A cube topped by a square pyramid (5 squares and 4 triangles): flatness is more evenly distributed. It requires $T=7$ tetrahedra, with $S_{i}=1$ and $N_{i}=7$.
Despite having the same surface flatness $S$, the second polyhedron requires more tetrahedra and internal gluing.
Finally, consider a polyhedron with many quadrilateral faces spread across all directions, such as the cube tunnel with $V=12$, $S=10$, and no external triangles. This structure also requires $T=T_{\min}+1$, even though surface flatness is maximal and triangle count is zero. The internal complexity in this case emerges from geometric necessity demonstrating that distributed flatness induces lettucing, even under strict constraints.

### v6

@@ 5.4 Count of Valid Internal Configurations
Empirical surface-type filtering at fixed $V$ (small-$V$ compression).
The ladder identities determine the full algebraic family $T=V-3+S_{i}$ and $N_{i}=V-4+2S_{i}$ under SALT+MIE. However, the realizability of the upper rungs depends on the external surface type $(E,F,S)$ at fixed $V$. Table LABEL:tab:V8-filter (compiled for $V=8$) exhibits a sharp compression: as external flatness $S$ increases (equivalently, as $E$ decreases so that E+S=3V6E+S=3V-6 remains fixed), progressively fewer $(T,N_{i},S_{i})$ rungs are realized.
In particular, for $S=6$ (the cube), only the two configurations with the lowest tetrahedral counts occur: $T=5$ (the $S_{i}=0$ rung) and $T=6$ (the $S_{i}=1$ rung). The higher algebraically admissible rungs $T=7$ and $T=8$ do not appear, despite being allowed by the ladder identities. Thus, the observed maximum is $T=6$, not the algebraic ceiling $T=8$.
This suggests that at large $S$, internal gluing structure may vary substantially in planarity, allowing non-coplanar interna gluingl triangles $N_{i}$ while maintaining minimal internal segmentation $S_{i}$, but that this flexibility is strongly constrained in small polyhedra. For small $V$, the interaction between surface structure and layering constraints therefore appears to impose a much sharper upper bound on $T$ than is predicted by the global incidence identities alone.
@@ 5.5 Example for $V=8$
$E$ $S$ (External) $T$ (Tetrahedra) $N_{i}$ (Triangles) $S_{i}$ (Segments)
18 0 5 4 0
18 0 6 6 1
18 0 7 8 2
18 0 8 10 3
17 1 5 4 0
17 1 6 6 1
17 1 7 8 2
16 2 5 4 0
16 2 6 6 1
15 3 5 4 0
15 3 6 6 1
14 4 5 4 0
14 4 6 6 1
13 5 5 4 0
13 5 6 6 1
12 6 5 4 0
12 6 6 6 1
Note. Here $E$ denotes the number of boundary edges prior to triangulating non-triangular faces. For fixed $V=8$, the triangulated boundary satisfies E+S=3V6=18E+S=3V-6=18.
From these configurations, the following empirical tendencies are observed for eight-vertex polyhedra:
  • When $S=0$: all V4V-4 ladder configurations are realized.
  • When $S=1$: the top ladder rung $(S_{i}=S_{i,\max})$ is not realized.
  • When $2\leq S\leq S_{i,\max}$: only the two lowest ladder rungs $(S_{i}=0,1)$ are realized.
Surface flatness $S$ thus correlates with increased constraint on internal decomposition at small $V$, but it does not act as an independent control parameter. Instead, realizability appears to be governed by the necessary packing of internal tetrahedra under fixed $(V,N)$ constraints. For example, even with a fully triangulated surface ($S=6$), the cube admits only $T=5$ or $T=6$ tetrahedra under SALT+MIE.
These behaviors are likely to become increasingly complex as $V$ increases, due to the geometric and combinatorial expansion of allowable internal connectivity. Accordingly, while empirical tendencies can be identified at fixed small $V$ (e.g., $4\leq V\leq 8$), extrapolating them to larger polyhedra is likely to be misleading, as it obscures emergent phenomena such as multilayer volumetric structure.
5.6 Determining admissible tetrahedral ranges from polygonal surface data
We consider the following problem: given a closed genus–0 polyhedral surface specified by a multiset of polygonal faces, determine the admissible range of tetrahedral decompositions consistent with purely symbolic (incidence-based) constraints. Throughout, we restrict attention to normal–form decompositions: all vertices lie on the boundary, no interior vertices are introduced, and all tetrahedra are glued face-to-face. When we invoke SALT+MIE, we mean the further restricted construction class used in §4 (shell-aligned layering; no edge bisections; at most one new interior segment per shell layer).
5.6.1 Input data and boundary normalization
Let the surface consist of faces with degrees $\deg f\geq 3$. From this data we compute:
$V=\text{number of vertices},\qquad E=\text{number of boundary edges},\qquad F=\text{number of faces}.$
Define the surface flatness parameter
$S:=\sum_{f}(\deg f-3),$
which counts the number of triangulation segments required to triangulate the surface without introducing new vertices.
After triangulation, the boundary data become
$N:=F+S\quad\text{(boundary triangles)},\qquad E_{\partial}:=E+S\quad\text{(boundary edges)}.$
For genus–0 surfaces these satisfy
$N=2V-4,\qquad E_{\partial}=3V-6,$
so once $V$ is fixed the triangulated boundary counts $(N,E_{\partial})$ are fixed.
5.6.2 Interior variables and the SALT+MIE identification
Let
[…]
Here $S_{i}$ counts the interior edges needed to fully triangulate the union of internal gluing surfaces. In the SALT+MIE class used in §4, every interior edge arises as such a triangulation segment, so $S_{i}$ is also the total count of interior edges in the tetrahedral complex. We therefore use $S_{i}$ as the interior-edge variable throughout (and do not introduce a separate $E_{i}$).
5.6.3 Exact incidence identities (normal form)
The following identities hold for normal-form tetrahedralizations with triangulated boundary and no interior vertices.
Triangle incidence.
Each tetrahedron contributes four triangular faces; boundary triangles are used once and internal triangles are shared by two tetrahedra. Hence
$4T=N+2N_{i}.$ (1)
Euler characteristic.
Let $E_{\partial}$ denote boundary edges after triangulation, and let $S_{i}$ denote interior edges (equivalently: interior triangulation segments in SALT+MIE). The Euler characteristic of the 3–ball gives
$V-(E_{\partial}+S_{i})+(N+N_{i})-T=1.$ (2)
Substituting the genus–0 boundary relations N=2V4N=2V-4 and $E_{\partial}=3V-6$, (2) simplifies to the exact identity
$\boxed{T-N_{i}+S_{i}=1.}$ (3)
5.6.4 Solution family and feasibility constraints
Equations (1) and (3) parameterize all admissible triples $(T,N_{i},S_{i})$ by the single integer $T$. Solving (1) for $N_{i}$ and substituting into (3) yields
$\boxed{N_{i}(T)=2T-\frac{N}{2},\qquad S_{i}(T)=N_{i}(T)-T+1=T-\frac{N}{2}+1.}$ (4)
Thus, for a fixed polygonal surface (hence fixed $V$ and fixed N=2V4N=2V-4), all admissible tetrahedralizations lie on the one-parameter integer family (4), subject only to nonnegativity:
$N_{i}(T)\geq 0,\qquad S_{i}(T)\geq 0.$
From $N_{i}(T)\geq 0$ one obtains the general incidence lower bound
$T\geq\left\lceil\frac{N}{4}\right\rceil=\left\lceil\frac{2V-4}{4}\right\rceil=\left\lceil\frac{V-2}{2}\right\rceil,$
which is necessary in any normal-form tetrahedralization. In the restricted SALT+MIE class, the stronger construction minimum $T_{\min}=V-3$ is observed (§4), and the ladder family below realizes all intermediate values.
5.6.5 Restricted SALT+MIE ranges (construction-based)
Within the SALT+MIE framework of §5 (at most one new interior segment per layer, and the first/last layers cannot host a new segment), one has
$0\leq S_{i}\leq V-5.$
Substituting into the ladder relations of §4 gives the construction-based ranges
$T$ $=V-3+S_{i},$ (5)
$N_{i}$ $=V-4+2S_{i},$ (6)
hence
$T_{\min}$ $=V-3,$ $T_{\max}$ $=2(V-4),$ (7)
$N_{i,\min}$ $=V-4,$ $N_{i,\max}$ $=3(V-4)-2,$ (8)
$S_{i,\max}$ $=V-5.$ (9)
These SALT+MIE bounds are not independent: they are linked by the exact identity (3).
Remark (Toward realizability criteria beyond the SALT ladder).
The results above distinguish three layers of constraint on internal tetrahedralization: exact incidence identities, construction-based restrictions (SALT+MIE), and empirical realizability observed in small cases. The first two layers generate a complete algebraic ladder of admissible $(T,N_{i},S_{i})$ configurations at fixed $V$, but they do not determine which rungs are realized for a given surface type $(E,F,S)$.
The $V=8$ examples show that this gap is substantive rather than technical. Although all ladder rungs are permitted algebraically, realizability depends strongly on how surface flatness is distributed across the boundary. For small polyhedra this produces sharp truncations of the ladder, while for larger $V$ similar constraints may degrade gradually, giving rise to multilayer interior structure rather than hard cutoffs. This is yet to be explored.